# Multimodal Unlearnable Examples: Protecting Data against Multimodal Contrastive Learning

### v1

@@ 4.4. Case Study: Face-Privacy Protection
Here, we conduct a case study to apply our MEM noise to a real-world scenario: protecting personal face images and associated information on social media platforms, such as names. This scenario is arguably one of the most common multimodal privacy protection scenarios. In previous sections, we assumed that hackers train models with multi-modal unlearnable examples from scratch. However, in most real-world scenarios, hackers will opt to fine-tune pre-trained models provided by the community. Therefore, here we assume that the protector aims to prevent their facial images and names from being fine-tuned with a pre-trained CLIP model. We adopt the Open AI’s released parameters to initialize our pre-trained models. We still assume the protector has access to their image and text data to generate unlearnable examples before sharing them on on-line social media platforms. However, they are unable to know the hacker’s model architecture and the parameters of pre-trained models. In this black-box scenario, we explore whether MEM can still prevent model from learning the protector’s face images.
We conducted experiments using the PubFig database (Kumar et al., 2009), a large real-world face dataset comprising 58,797 images of 200 individuals collected from the internet. For retrieval evaluation, we randomly selected one photo of each celebrity as the test set and utilized all remaining images for training. It’s noteworthy that the pre-trained CLIP model already possesses knowledge of the real names and faces of these celebrities. Therefore, for authentic fine-tuning, we altered their names and provided a set of text templates relevant to the name for caption generation. Subsequently, we generated unlearnable examples using MEM and employed a pre-trained surrogate model with ResNet50 as the image encoder. We fine-tune the pre-trained models for 10 epochs and a small learning rate of $0.00001$ and evaluate them with different hacker models. The results are presented in Table 4. For comparison, we also test the results of the pre-trained models and the benign fine-tuned models. Table 4 shows that the initial model shows poor performance on the test set, since the face and name features are not aligned. However, with only 10 epochs of fine-tuning, the Medr of its retrieved results reaches 1. This may be because of the strong feature representation of the pre-trained model, enabling it to swiftly adapt to the new task during the fine-tuning process. However, MEM can prevent these fine-tuned models from learning the correlation between face and name features, thereby impeding accurate person retrieval on the test set. Our unlearnable examples were generated by ResNet50 surrogate model and effectively protect on the ResNet101 hacker model. However, their efficacy is considerably reduced on ViT, possibly due to variations in architecture and initialization parameters.

### v2

@@ 4.4. Case Study: Face-Privacy Protection
We conduct a case study to apply MEM to a real-world scenario: protecting personal face images and associated information on social media platforms, such as names. It is arguably one of the most common multimodal privacy protection scenarios. In previous parts, we assumed that hackers train models with unlearnable examples from scratch, but in the real-world, hackers will opt to fine-tune pre-trained models provided by the community. Therefore, here we assume that the protector aims to prevent their facial images and names from being fine-tuned with a pre-trained CLIP model. We adopt the Open AI’s released parameters to initialize pre-trained models, and still assume the protector has access to their data to generate unlearnable examples before sharing them on on-line platforms. However, they are unable to know the hacker’s model architecture and the parameters. In this black-box scenario, we explore whether MEM can still prevent model from learning.
We conducted experiments using the PubFig  (Kumar et al., 2009), a large real-world face dataset comprising 58,797 images of 200 individuals. For retrieval evaluation, we randomly selected one photo of each celebrity as the test set and utilized remaining images for training. It’s noteworthy that the pre-trained CLIP model already possesses knowledge of the real names and faces of these celebrities. For authentic fine-tuning, we altered their names and provided a set of text templates for caption generation, as shown in Fig. 7. Then we generated unlearnable examples using MEM and employed a pre-trained surrogate model with ResNet50. We fine-tune the pre-trained models for 10 epochs and a small learning rate of $0.00001$ and evaluate them with different hacker models. The results are presented in Tab. 4. For comparison, we also test the results of the pre-trained models and the benign fine-tuned models. We find the initial model shows poor performance on the test set, since the face and name features are not aligned. However, with only 10 epochs of fine-tuning, the Medr of its retrieved results reaches 1. This could be because the pre-trained model has a strong feature representation, enabling it to rapidly adapt to the new task in the fine-tuning process. Our MEM can prevent these models from learning the correlation between face and name features, thereby impeding accurate person retrieval on the test set. These unlearnable examples were generated by ResNet50 surrogate model and effectively work on the ResNet101 targeted model. However, their efficacy is reduced on ViT, possibly due to variations in architecture and initialization parameters.
Figure 7. Illustration of the face-privacy protection pipeline.